# Sufficient conditions for QMC analysis of finite elements for

### v1

@@ Abstract
Parametric regularity of discretizations of flux vector fields satisfying a balance law is studied under some assumptions on a random parameter that links the flux with an unknown primal variable (often through a constitutive law). In the primary example of the stationary diffusion equation, the parameter corresponds to the inverse of the diffusivity. The random parameter is modeled here as a Gevrey-regular random field. Specific focus is on random fields expressible as functions of countably infinite sequences of independent random variables, which may be uniformly or normally distributed. Quasi-Monte Carlo (QMC) error bounds for some quantity of interest that depends on the flux are then derived using the parametric regularity. It is shown that the QMC method converges optimally if the quantity of interest depends continuously on the primal variable, its flux, or its gradient. A series of assumptions are introduced with the goal of encompassing a broad class of discretizations by various finite element methods. The assumptions are verified for the diffusion equation discretized using conforming finite elements, mixed methods, and hybridizable discontinuous Galerkin schemes. Numerical experiments confirm the analytical findings, highlighting the role of accurate flux approximation in QMC methods.
@@ 1 Introduction
Under further assumptions, we establish dimension-independent error bounds for quasi-Monte Carlo (QMC) methods applied to compute expected values of quantities of interest. These quantities are assumed to depend continuously on either the solution $u$, the flux $\boldsymbol{q}$, or the gradient $\nabla u$. Our analysis shows that optimal convergence rates of QMC methods can be achieved in this setting, provided the integrands inherit sufficient regularity from the underlying parametric dependence.
@@ 1.1 The prominent example of the diffusion equation
The parametric diffusion equation solves for a function $u$ of parameter $\omega$ and spatial variable $\boldsymbol{x}$. Here $\omega$ is a random parameter sample from a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and $\boldsymbol{x}$ is an element of the spatial domain $D\subset\mathbb{R}^{d}$, a bounded open set generally of much smaller dimension (e.g., $d\in\{1,2,3\}$) than the parameter space. Let $L^{2}(D)$ denotes the space of square-integrable functions on $D$ and let $H^{1}_{0}(D)$ denote its standard Sobolev subspace of functions having square-integrable first-order weak derivatives and vanishing trace on the spatial boundary $\partial D$. The parameter-dependent solution $u\in L^{2}(\Omega,\mathcal{F},\mathbb{P};H^{1}_{0}(D))$ is such that for almost every $\omega\in\Omega$,
@@
where $f\in L^{2}(D)$, and $a(\omega),a^{-1}(\omega)\in L^{\infty}(D)$ are positive functions on $D$. In many applications, the coefficient $a$ is used to model spatially varying material properties, with uncertainties described probabilistically (cf., e.g., [6]). We follow the common approach where $a^{-1}$ is assumed to be derived from a random field, i.e., we identify
@@
where $\boldsymbol{y}=(y_{m})_{m\in\mathbb{N}}\in U\subset\mathbb{R}^{\mathbb{N}}$ is a sequence of independent random variables on the probability space $(\Omega,\mathcal{F},\mathbb{P})$. Examples of $U$ are given in Section 2. In this work, we consider a problem in which the reciprocal of the random field $a$ enters (1)—although this is not a significant change in the lognormal setting. This setting leads naturally to mixed finite element methods that are exceptionally useful if the quantity of interest is related to the flux of the flow field of some physical phenomenon.
@@ 1.2 Formulations in terms of fluxes
In particular, mixed methods rewrite (1) as two subproblems, namely, (i) the balance equation $\nabla\cdot\boldsymbol{q}(\boldsymbol{y})=f$, which is the diffusion equation if (ii) the constitutive equation is modeled according to Fick’s law $\boldsymbol{q}(\boldsymbol{y})=-a^{-1}(\boldsymbol{y})\nabla u(\boldsymbol{y})$. This manuscript investigates a unified framework that allows the analysis of parametric regularity bounds of such continuity equations for several finite element schemes using the parametric diffusion equation (1) as a fundamental example.
While many authors have already investigated the (continuous, classic) finite element method for (1), see the references in Section 1.3, there are very few works exploiting the properties of the dual variable $\boldsymbol{q}(\boldsymbol{y})$. Notable examples thereof include [20] that uses multilevel Monte Carlo methods, [18] that applies mixed methods in concert with quasi-Monte Carlo (QMC) methods, and [14], which considers the mixed formulation for lognormal random fields, for which $a^{-1}$ is also lognormal. The latter proves analytic dependence on the parameter in this case, yielding a parametric regularity result for the undiscretized flux.
@@ 1.3 Parametrization of randomness
The majority of studies on the application of QMC methods to PDEs with random coefficients have been carried out under the assumption of a fixed parametric representation for the input random field: some commonly used parametric models include the so-called uniform and affine model [7, 11, 13, 15, 40, 31, 32, 39] and the lognormal model [16, 17, 18, 19, 24, 30, 41]. Recently, several studies have gone beyond these models, focusing instead on the class of Gevrey regular input random fields [3, 4, 22]. This class contains infinitely smooth but generally non-analytic functions with a growth condition imposed on their higher-order partial derivatives. Such representations of input uncertainty can be significantly more general than those previously considered in the literature, and it has been demonstrated in [3, 4, 22] that it is possible to construct QMC cubature rules that achieve dimension-independent, faster-than-Monte Carlo cubature convergence rates for the quantification of uncertainties in stochastic models. This has also been done under abstract/general assumptions to the considered problem [21].
In particular, the works above show that if the input random field is sufficiently smooth with respect to the uncertain variables, then the PDE response will also be smooth. The benefit of using cubature methods such as QMC over Monte Carlo methods is that they can exploit the smoothness of the PDE response, yielding faster convergence rates for the computation of the response statistics. Specific QMC integration methods—e.g., lattice rules [29, 29, 12, 10]—can be trivially implemented in parallel, making them ideal for large-scale computations.
@@ 1.4 Notation
Infinite multi-indices are sequences of non-negative integers of the form ${\boldsymbol{{\nu}}}=(\nu_{1},\nu_{2},\dots)$. Their length, defined by $|\boldsymbol{\nu}|:=\sum_{j=1}^{\infty}\nu_{j}$, is generally infinite. Let $\mathscr{F}$ denote the set of infinite multi-indices $\boldsymbol{{\nu}}$} of finite length $|\boldsymbol{\nu}|<\infty$, i.e., $\mathscr{F}$ is the set of all finitely supported infinite multi-indices. Also let
@@ 2 Quasi-Monte Carlo method
The quasi-Monte Carlo (QMC) method recalled in this section can approximate integrals of functions supported in a high-dimensional cube. Let $F\colon\mathcal{I}^{s}\to\mathbb{R}$ be a continuous function, where $s\in\mathbb{N}$ denotes some (high) dimension and $\mathcal{I}$ denotes an interval (not necessarily bounded) of $\mathbb{R}$. Given a nonzero probability density function $\varphi:\mathcal{I}\to(0,\infty)$, we are interested in approximating the integral
@@
i.e., the integral of $F$ with respect to the products of the measures generated by $\varphi$ on $\mathcal{{I}}$}, assuming it exists.
@@ 2.1 Bounded domain $\tfrac{1}{2},\tfrac{1}{2}]^{s}$
Consider the case of $\mathcal{I}^{s}=[-\tfrac{1}{2},\tfrac{1}{2}]$. Since the Lebesgue measure of $\tfrac{1}{2},\tfrac{1}{2}]^{s}$ is one, its indicator function
@@
where the sum runs over all subsets $\mathfrak{u}$ of $\{1:s\}$, the inner integral (${\rm d}\boldsymbol{y}_{-\mathfrak{u}}$) is over all $y_{j}$ with $j\in\{1:s\}\setminus\mathfrak{u}$ (called the inactive variables), while the outer integral (${\rm d}\boldsymbol{y}_{\mathfrak{u}}$) is over all $y_{j}$ with $j\in\mathfrak{u}$ (called the active variables). It is known [8, 12, 36] that a generating vector can be obtained by a component-by-component (CBC) algorithm resulting in a QMC rule with the following error bound.
Theorem 2.1.
@@ 2.2 Unbounded domain $\mathbb{R}^{s}$
Then $\hat{\phi}^{-1}(t)=\frac{1}{2}\big(1+{\rm erf}\big(\frac{t}{\sqrt{2}}\big)\big)$ for $t\in(0,1)$ where ${\rm erf}(\cdot)$ denotes the Gauss error function. This $\varphi$ corresponds to the probability density function of the multivariate normal distribution $\mathcal{N}(\mathbf{0},\mathbf{I}_{s\times s})$ of vanishing mean and variance $\mathbf{I}_{s\times s}$, the $s\times s$ identity matrix.
To describe the regularity needed from $F$ in this case, suppose we are given a $\gamma$ as before (see Section 2.1). In addition, suppose we are given a sequence of positive numbers $\alpha=(\alpha_{j}),j\in{\mathbb{{N}}}.$ Letting
@@
Then it is known that a generating vector can be obtained by a CBC algorithm [35] satisfying the following rigorous error bound.
Theorem 2.2 (cf. [17, Theorem 15]).
@@ 3.1 The setting
We want to approximate functionals of a flux $\boldsymbol{{q}}$} that depend on some random coefficient $a({\boldsymbol{{x}}},\omega)$ where $\boldsymbol{{x}}$} is in $D$, a Lipschitz polyhedron in ${\mathbb{{R}}}^{d}$, and $\omega$ is a sample point in a probability space $(\Omega,\mathcal{F},\mathbb{P})$. Following prior authors (see e.g., [29]) we assume that the random coefficient $a({\boldsymbol{{x}}},\omega)$ has been parameterized by a vector ${\boldsymbol{{y}}}=(y_{1}(\omega),y_{2}(\omega),\cdots)$ where every component $y_{j}$ lies in some (bounded or unbounded) real interval $\mathcal{{I}}$}, i.e., $\boldsymbol{{y}}$} is in $U={\mathcal{{I}}}^{{\mathbb{N}}}$. Endowing $\mathcal{{I}}$} with a probability measure generated by a probability density function $\varphi$ on $\mathcal{{I}}$}, the set $U$ inherits a product measure which we denote by $\mu(d{\boldsymbol{{y}}})$. Writing $a({\boldsymbol{{y}}})$ for $a(\cdot,{\boldsymbol{{y}}})$, we assume that $a({\boldsymbol{{y}}})>0$ a.e. in $D$ and that $a$ is Gevrey-$\sigma$ regular, a property that will be rigorously formulated in (A5).
The exact flux $\boldsymbol{q}$ is in $L^{2}(\Omega,\mathcal{F},\mathbb{P};L^{2}(D))$ and has the property that $\boldsymbol{q}(\omega)\in H(\operatorname{div},D)$ for almost every $\omega\in\Omega$. While $\boldsymbol{{q}}$} solves the balance law $\operatorname{div}{\boldsymbol{{q}}}=f$, its discrete approximation ${\boldsymbol{{q}}}_{h}$ in some (conforming or nonconforming) finite element space $Q_{h}$ with norm $\|\cdot\|_{Q_{h}}$, solves a discretization of the law (such as those considered in Section 4). Here $h$ denotes some meshsize parameter related to the discretization. The dependence of the exact and discrete fluxes on the random coefficient is henceforth indicated by ${\boldsymbol{{q}}}({\boldsymbol{{y}}})$ and ${\boldsymbol{{q}}}_{h}({\boldsymbol{{y}}})$, which are both functions of the spatial variable $\boldsymbol{{x}}$} in $D$, and are in $H(\operatorname{div},D)$ and $Q_{h}$, respectively for $\boldsymbol{{y}}$} in $U$.
Suppose a quantity of interest is represented by a real-valued (not necessarily linear) functional $J$ of the flux, $J:H(\operatorname{div},D)+Q_{h}\to{\mathbb{{R}}}$, also called the goal functional. A typical example is $J({\boldsymbol{{q}}})=\int_{D}a^{-1}{\boldsymbol{{q}}}\cdot{\boldsymbol{{q}}}\;\;\textup{d}\boldsymbol{x}$ (see Example 3.2 below). We are interested in approximating the expected value of $J({\boldsymbol{{q}}}({\boldsymbol{{y}}}))$ with respect to the random variables parameterized by $\boldsymbol{{y}}$}, namely
@@
where the integral is computed using the above-mentioned product measure $\mu(\textup{d}{\boldsymbol{{y}}})$ generated by $\varphi$. To approximate this using the QMC rule of Section 2, we first truncate the infinite sequence $\boldsymbol{{y}}$} to its first $s$ components ${\boldsymbol{{y}}}_{s}={\boldsymbol{{q}}}(y_{1},\dots,y_{s},0,\dots)$ and set ${\boldsymbol{{q}}}_{s}={\boldsymbol{{q}}}({\boldsymbol{{y}}}_{s})$. A computable finite element flux also employs the same $s$-components, ${\boldsymbol{{q}}}_{h}\equiv{\boldsymbol{{q}}}_{h}({\boldsymbol{{y}}}_{s})$. Having truncated the parameter space dimension to the finite dimension $s$, we may now use the QMC rule of Section 2 to approximate the integral. Summarizing these standard steps, the final approximation error is split into dimension truncation error, finite element error, and cubature error, namely
@@
We now state a few conditions under which the last term representing the QMC error after dimension truncation can be bounded. Henceforth, we assume that $\boldsymbol{{y}}$} is already dimension-truncated, i.e., ${\boldsymbol{{y}}}\equiv{\boldsymbol{{y}}}_{s}$ from now on. Note however that all bounds we derive will be $s$-independent, so the dimension $s$ can be made arbitrarily close to infinity without any deterioration in the constants. In our conditions below, all constants denoted by $C$ with some subscript are assumed to be independent of $s$ and $h$. The first two conditions below are usually immediate in methods generating flux approximations.
  1. (A1)
    $\|\cdot\|_{Q_{h}}$ is stronger than the energy norm. There is a constant $C_{E}>0$ such that
    $\left\|\sqrt{a(\boldsymbol{y})}\;\frac{\partial^{|\mathfrak{u}|}}{\partial\boldsymbol{y}_{\mathfrak{u}}}\boldsymbol{q}_{h}(\boldsymbol{y})\right\|_{L^{2}(D)}\leq C_{E}\left\|\frac{\partial^{|\mathfrak{u}|}}{\partial\boldsymbol{y}_{\mathfrak{u}}}\boldsymbol{q}_{h}(\boldsymbol{y})\right\|_{Q_{h}}$
    for all $\boldsymbol{y}\in U$ and all subsets $\mathfrak{u}$ of $\{1:s\}$.
  2. (A2)
    Stability of the numerical method. There is a constant $C_{S}>0$ such that
    $\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}\leq C_{S}\|a(\boldsymbol{y})\|^{1/2}_{L^{\infty}(D)}\|f\|_{L^{2}(D)}\qquad\text{ for all }\boldsymbol{y}\in U.$
@@
  1. (A3)
    Recursive bound for the parametric derivative. There is $C_{R}>0$ such that for any $\boldsymbol{\nu}\in\mathscr{F}_{1}$ (see (2)) and $\boldsymbol{y}\in U$, we have
    $\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}\leq C_{R}\sum_{\boldsymbol{0}\neq\boldsymbol{m}\leq\boldsymbol{\nu}}\begin{pmatrix}\boldsymbol{\nu}\\ \boldsymbol{m}\end{pmatrix}\left\|\frac{\partial^{\boldsymbol{m}}_{\boldsymbol{y}}a(\boldsymbol{y})}{a(\boldsymbol{y})}\right\|_{L^{\infty}(D)}\|\partial^{\boldsymbol{\nu}-\boldsymbol{m}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}.$
@@
  1. (A4)
    Boundedness of the quantity of interest. There is $C_{J}>0$ and $r>0$ such that for any $\mathfrak{u}\subset\{1:s\}$ and every $\boldsymbol{y}\in U$,

### v2

@@ Abstract
Parametric regularity of discretizations of flux vector fields satisfying a balance law is studied under some assumptions on a random parameter that links the flux with an unknown primal variable (often through a constitutive law). In the primary example of the stationary diffusion equation, the parameter corresponds to the inverse of the diffusivity. The random parameter is modeled here as a Gevrey-regular random field. Specific focus is on random fields expressible as functions of countably infinite sequences of independent random variables, which may be uniformly or normally distributed. Quasi-Monte Carlo (QMC) error bounds for some quantity of interest that depends on the flux are then derived using the parametric regularity. It is shown that the QMC method achieves a dimension-independent, faster-than-Monte Carlo convergence rate if the quantity of interest depends continuously on the primal variable, its flux, or its gradient. A series of assumptions are introduced with the goal of encompassing a broad class of discretizations by various finite element methods. The assumptions are verified for the diffusion equation discretized using conforming finite elements, mixed methods, and hybridizable discontinuous Galerkin schemes. Numerical experiments confirm the analytical findings, highlighting the role of accurate flux approximation in QMC methods.
@@ 1 Introduction
Under further assumptions, we establish dimension-independent error bounds for quasi-Monte Carlo (QMC) methods applied to compute expected values of quantities of interest. These quantities are assumed to depend continuously on either the solution $u$, the flux $\boldsymbol{q}$, or the gradient $\nabla u$. Our analysis shows that dimension-independent and faster-than-Monte Carlo convergence rates of QMC methods can be achieved in this setting, provided the integrands inherit sufficient regularity from the underlying parametric dependence.
@@ 1.1 The prominent example of the diffusion equation
The parametric diffusion equation solves for a function $u$ of parameter $\omega$ and spatial variable $\boldsymbol{x}$. Here $\omega$ is a random parameter sample from a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and $\boldsymbol{x}$ is an element of the spatial domain $D\subset\mathbb{R}^{d}$, a bounded open set with typically small spatial dimension (e.g., $d\in\{1,2,3\}$). Let $L^{2}(D)$ denotes the space of square-integrable functions on $D$ and let $H^{1}_{0}(D)$ denote its standard Sobolev subspace of functions having square-integrable first-order weak derivatives and vanishing trace on the spatial boundary $\partial D$. The parameter-dependent solution $u\in L^{2}(\Omega,\mathcal{F},\mathbb{P};H^{1}_{0}(D))$ is such that for almost every $\omega\in\Omega$,
@@
where $f\in L^{2}(D)$, and $a(\omega),a^{-1}(\omega)\in L^{\infty}(D)$ are positive functions on $D$. In many applications, the coefficient $a$ is used to model spatially varying material properties, with uncertainties described probabilistically (cf., e.g., [cdv]). We follow the common approach where $a^{-1}$ is assumed to be derived from a random field, i.e., we identify
@@
where $\boldsymbol{y}=(y_{m})_{m\in\mathbb{N}}\in U\subset\mathbb{R}^{\mathbb{N}}$ is a sequence of independent random variables on the probability space $(\Omega,\mathcal{F},\mathbb{P})$. Examples of $U$ are given in Section 2. In practice, the infinite-dimensional parameter set $U$ is truncated into an $s$-dimensional space $U_{s}$, where $s$ is the truncation dimension, not to be confused with the spatial dimension $d$. In this work, we consider a problem in which the reciprocal of the random field $a$ enters (1)—although this is not a significant change in the lognormal setting. This setting leads naturally to mixed finite element methods that are exceptionally useful if the quantity of interest is related to the flux of the flow field of some physical phenomenon.
@@ 1.2 Formulations in terms of fluxes
Virtually all analyses of quasi-Monte Carlo methods for uncertainty quantification of partial differential equations with random coefficients employ a spatially conforming, primal finite element discretization. However, in many cases, it is preferable to consider non-conforming or dual finite element methods. For example, in groundwater flow simulations and coupled (contaminant) transport problems, such schemes are essential for obtaining mass conservation. Our analysis provides an abstract framework in which primal, dual, and several other methods can be analyzed in the same manner if the inverse of the random coefficient is Gevrey regular, cf. Section 1.3.
Mixed formulations rewrite (1) as two subproblems, namely, (i) the balance equation $\nabla\cdot\boldsymbol{q}(\boldsymbol{y})=f$, which is the diffusion equation if (ii) the constitutive equation is modeled according to Fick’s law $\boldsymbol{q}(\boldsymbol{y})=-a^{-1}(\boldsymbol{y})\nabla u(\boldsymbol{y})$. This manuscript investigates a unified framework that allows the analysis of parametric regularity bounds of such continuity equations for several finite element schemes using the parametric diffusion equation (1) as a fundamental example.
While many authors have already investigated the (continuous, classic) finite element method for (1), see the references in Section 1.3, there are very few works exploiting the properties of the dual variable $\boldsymbol{q}(\boldsymbol{y})$. Notable examples thereof include [GrahamSU15] that uses multilevel Monte Carlo methods, [GrahamKNSS11] that applies mixed methods in concert with quasi-Monte Carlo (QMC) methods, and [ErnstS14], which considers the mixed formulation for lognormal random fields, for which $a^{-1}$ is also lognormal. The latter proves analytic dependence on the parameter in this case, yielding a parametric regularity result for the undiscretized flux.
@@ 1.3 Parametrization of randomness
The majority of studies on the application of QMC methods to PDEs with random coefficients have been carried out under the assumption of a fixed parametric representation for the input random field: some commonly used parametric models include the so-called uniform and affine model [cds10, spodpaper14, dicklegiaschwab, ghs18, SchwabG11, kss12, kssmultilevel, schwab13] and the lognormal model [gittelson, GrahamKNSSS15, GrahamKNSS11, log3, log4, log5, schwabtodor]. Recently, several studies have gone beyond these models, focusing instead on the class of Gevrey regular input random fields [ChernovL24, ChernovL24b, Harbrecht24]. This class contains infinitely smooth but generally non-analytic functions with a growth condition imposed on their higher-order partial derivatives. Such representations of input uncertainty can be significantly more general than those previously considered in the literature, and it has been demonstrated in [ChernovL24, ChernovL24b, Harbrecht24] that it is possible to construct QMC cubature rules that achieve dimension-independent, faster-than-Monte Carlo cubature convergence rates for the quantification of uncertainties in stochastic models. This has also been done under abstract/general assumptions to the considered problem [GuKa24].
In particular, the works above show that if the input random field is sufficiently smooth with respect to the uncertain variables, then the PDE response will also be smooth. The benefit of using cubature methods such as QMC over Monte Carlo methods is that they can exploit the smoothness of the PDE response, yielding faster convergence rates for the computation of the response statistics. Specific QMC integration methods—e.g., lattice rules [kuonuyenssurvey, kuonuyenssurvey, dks13, dick2022lattice]—can be trivially implemented in parallel, making them ideal for large-scale computations.
While there exists a theory for higher-order QMC in the uniform and bounded setting [spodpaper14, hoqmcpaper], there is currently no complete higher-order QMC theory in the unbounded setting. To this end, we focus on lattice rules, which can be shown to work in both settings. That is, the aim of this work is to develop a cohesive theoretical framework for both, the bounded and unbounded settings rather than extending the higher-order QMC theory.
@@ 1.4 Notation
Infinite multi-indices are sequences of non-negative integers of the form ${\boldsymbol{{\nu}}}=(\nu_{1},\nu_{2},\dots)$. Let $\mathscr{F}$ denote the set of infinite multi-indices $\boldsymbol{{\nu}}$} of finite length $|\boldsymbol{\nu}|{:=\sum_{j=1}^{\infty}\nu_{j}}<\infty$, i.e., $\mathscr{F}$ is the set of all finitely supported infinite multi-indices. Also let
@@ 2 Quasi-Monte Carlo method
The quasi-Monte Carlo (QMC) method recalled in this section can approximate integrals of functions supported in a high-dimensional cube. Let $F\colon\mathcal{I}^{s}\to\mathbb{R}$ be a continuous function, where $s\in\mathbb{N}$ denotes some (high) dimension and $\mathcal{I}$ denotes an interval (not necessarily bounded) of $\mathbb{R}$. Given a strictly positive probability density function $\varphi:\mathcal{I}\to(0,\infty)$, we are interested in approximating the integral
@@
i.e., the integral of $F$ with respect to the product measure on $\mathcal{I}^{s}$ induced by the density $\varphi$ on $\mathcal{{I}}$}, provided that $F$ is integrable.
@@ 2.1 Bounded domain $\tfrac{1}{2},\tfrac{1}{2}]^{s}$
Consider the case of $\mathcal{I}^{s}=[-\tfrac{1}{2},\tfrac{1}{2}]^{s}$. Since the Lebesgue measure of $\tfrac{1}{2},\tfrac{1}{2}]^{s}$ is one, its indicator function
@@
where the sum runs over all subsets $\mathfrak{u}$ of $\{1:s\}$, the inner integral (${\rm d}\boldsymbol{y}_{-\mathfrak{u}}$) is over all $y_{j}$ with $j\in\{1:s\}\setminus\mathfrak{u}$ (called the inactive variables), while the outer integral (${\rm d}\boldsymbol{y}_{\mathfrak{u}}$) is over all $y_{j}$ with $j\in\mathfrak{u}$ (called the active variables). It is known [ckn06, dks13, cn06] that a generating vector can be obtained by a component-by-component (CBC) algorithm resulting in a QMC rule with the following error bound.
Theorem 2.1 ([kuonuyenssurvey, Theorem 5.1]).
@@ 2.2 Unbounded domain $\mathbb{R}^{s}$
Then $\hat{\phi}^{-1}(t)=\frac{1}{2}\big(1+{\rm erf}\big(\frac{t}{\sqrt{2}}\big)\big)$ for $t\in(0,1)$ where ${\rm erf}(\cdot)$ denotes the Gauss error function. This $\varphi$ corresponds to the probability density function of the multivariate normal distribution $\mathcal{N}(\mathbf{0},\mathbf{I}_{s\times s})$ of vanishing mean and variance $\mathbf{I}_{s\times s}$, the $s\times s$ identity matrix. Notably, using this specific function is a choice that we make for our manuscript. This choice, however, is not canonical and is possibly not ideal as indicated in [KazashiSG25, KritzerPPW20].
To describe the regularity needed from $F$ in this case, suppose we are given a $\gamma$ as before (see Section 2.1). In addition, suppose we are given a sequence of positive numbers $\alpha=(\alpha_{j})_{{j\in{\mathbb{{N}}}}}.$ Letting
@@
Then it is known that a generating vector can be obtained by a CBC algorithm [nicholskuo] satisfying the following rigorous error bound.
Theorem 2.2 (cf. [nicholskuo, Theorem 7]).
@@ 3.1 The setting
We want to approximate functionals of a flux $\boldsymbol{{q}}$} that depend on some random coefficient $a({\boldsymbol{{x}}},\omega)$ where $\boldsymbol{{x}}$} is in $D$, a Lipschitz polyhedron in ${\mathbb{{R}}}^{d}$, and $\omega$ is a sample point in a probability space $(\Omega,\mathcal{F},\mathbb{P})$. Following prior authors (see e.g., [kuonuyenssurvey]) we assume that the random coefficient $a({\boldsymbol{{x}}},\omega)$ has been parameterized by a vector ${\boldsymbol{{y}}}=(y_{1}(\omega),y_{2}(\omega),\cdots)$ where every component $y_{j}$ lies in some (bounded or unbounded) real interval $\mathcal{{I}}$}, i.e., $\boldsymbol{{y}}$} is in $U={\mathcal{{I}}}^{{\mathbb{N}}}$. Endowing $\mathcal{{I}}$} with a probability measure generated by a probability density function $\varphi$ on $\mathcal{{I}}$}, the set $U$ inherits a product measure which we denote by $\mu({\rm d}{\boldsymbol{{y}}})$. Writing $a({\boldsymbol{{y}}})$ for $a(\cdot,{\boldsymbol{{y}}})$, we assume that $a({\boldsymbol{{y}}})>0$ a.e. in $D$ (i.e., for each $\boldsymbol{y}$ we have $a(\boldsymbol{y})=a(\boldsymbol{x},\boldsymbol{y})>0$ for Lebesgue-a.e. $\boldsymbol{x}\in D$) and that $a$ is Gevrey-$\sigma$ regular, a property that will be rigorously formulated in 5.
The exact flux $\boldsymbol{q}$ is in $L^{2}(\Omega,\mathcal{F},\mathbb{P};L^{2}(D))$ and has the property that $\boldsymbol{q}(\omega)\in H(\operatorname{div},D)$ for almost every $\omega\in\Omega$. While $\boldsymbol{{q}}$} solves the balance law $\operatorname{div}{\boldsymbol{{q}}}=f$, its discrete approximation ${\boldsymbol{{q}}}_{h}$ in some (conforming or nonconforming) finite element space $\mathcal{W}_{h}$ with norm $\|\cdot\|_{\mathcal{W}_{h}}$, solves a discretization of the law (such as those considered in Section 4). Here $h$ denotes some meshsize parameter related to the discretization. The dependence of the exact and discrete fluxes on the random coefficient is henceforth indicated by ${\boldsymbol{{q}}}({\boldsymbol{{y}}})$ and ${\boldsymbol{{q}}}_{h}({\boldsymbol{{y}}})$, which are both functions of the spatial variable $\boldsymbol{{x}}$} in $D$, and are in $H(\operatorname{div},D)$ and $\mathcal{W}_{h}$, respectively for $\boldsymbol{{y}}$} in $U$.
Suppose a quantity of interest is represented by a real-valued (not necessarily linear) functional $J$ of the flux, $J:H(\operatorname{div},D)+\mathcal{W}_{h}\to{\mathbb{{R}}}$, also called the goal functional. A typical example is $J({\boldsymbol{{q}}})=\int_{D}a^{-1}{\boldsymbol{{q}}}\cdot{\boldsymbol{{q}}}\;\;\textup{d}\boldsymbol{x}$ (see Example 3.2 below). We are interested in approximating the expected value of $J({\boldsymbol{{q}}}({\boldsymbol{{y}}}))$ with respect to the random variables parameterized by $\boldsymbol{{y}}$}, namely
@@
where the integral is computed using the above-mentioned product measure $\mu(\textup{d}{\boldsymbol{{y}}})$ generated by $\varphi$. To approximate this using the QMC rule of Section 2, we first truncate the infinite sequence $\boldsymbol{{y}}$} to its first $s$ components ${\boldsymbol{{y}}}_{s}=(y_{1},\dots,y_{s},0,0,\dots){\in U}$ and set ${\boldsymbol{{q}}}_{s}={\boldsymbol{{q}}}({\boldsymbol{{y}}}_{s})$. A computable finite element flux also employs the same $s$ components, ${\boldsymbol{{q}}}_{h}\equiv{\boldsymbol{{q}}}_{h}({\boldsymbol{{y}}}_{s})$. Where convenient, we occasionally abuse notation and identify $\boldsymbol{q}_{h}(y_{1},\ldots,y_{s})\equiv\boldsymbol{q}_{h}(y_{1},\ldots,y_{s},0,0,\ldots)$ for $\boldsymbol{y}\in U_{s}$. Having truncated the parameter space dimension to the finite dimension $s$, we may now use the QMC rule of Section 2 to approximate the integral. Summarizing these standard steps, the final approximation error is split into dimension truncation error, finite element error, and cubature error, namely
@@
We now state a few conditions under which the last term representing the QMC error after dimension truncation can be bounded. Henceforth, unless otherwise stated, we assume that $\boldsymbol{{y}}$} is already dimension-truncated, i.e., ${\boldsymbol{{y}}}\equiv{\boldsymbol{{y}}}_{s}$ from now on. Note however that all bounds we derive will be $s$-independent, so the dimension $s$ can be made arbitrarily close to infinity without any deterioration in the constants. In our conditions below, all constants denoted by $C$ with some subscript are assumed to be independent of $s$ and $h$. The first two conditions below are usually immediate in methods generating flux approximations. It refers to a general norm $\|\cdot\|_{\mathcal{W}_{h}}$ that can be chosen freely such that all assumptions hold in that norm.
Assumption 1.
$\|\cdot\|_{\mathcal{W}_{h}}$ is stronger than the energy norm. There is a constant $C_{E}>0$ such that for all $\boldsymbol{y}\in{U_{s}}$ and all subsets $\mathfrak{u}$ of $\{1:s\}$
$\left\|\sqrt{a(\boldsymbol{y})}\;\frac{\partial^{|\mathfrak{u}|}}{\partial\boldsymbol{y}_{\mathfrak{u}}}\boldsymbol{q}_{h}(\boldsymbol{y})\right\|_{L^{2}(D)}\leq C_{E}\left\|\frac{\partial^{|\mathfrak{u}|}}{\partial\boldsymbol{y}_{\mathfrak{u}}}\boldsymbol{q}_{h}(\boldsymbol{y})\right\|_{\mathcal{W}_{h}}.$
Assumption 2.
Stability of the numerical method. There is a constant $C_{S}>0$ with
$\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}\leq C_{S}\|a(\boldsymbol{y})\|^{1/2}_{L^{\infty}(D)}\|f\|_{L^{2}(D)}\qquad\text{ for all }\boldsymbol{y}\in{U_{s}}.$
@@
Assumption 3.
Recursive bound for the parametric derivative. There is $C_{R}>0$ such that for any $\boldsymbol{\nu}\in\mathscr{F}_{1}$ (see (2)) and $\boldsymbol{y}\in{U_{s}}$, we have
$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}\leq C_{R}\sum_{\boldsymbol{0}\neq\boldsymbol{m}\leq\boldsymbol{\nu}}\begin{pmatrix}\boldsymbol{\nu}\\ \boldsymbol{m}\end{pmatrix}\left\|\frac{\partial^{\boldsymbol{m}}_{\boldsymbol{y}}a(\boldsymbol{y})}{a(\boldsymbol{y})}\right\|_{L^{\infty}(D)}\|\partial^{\boldsymbol{\nu}-\boldsymbol{m}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}.$
@@
Assumption 4.
Boundedness of the quantity of interest. There is $C_{J}>0$ and $r>0$ such that for any $\mathfrak{u}\subset\{1:s\}$ and every $\boldsymbol{y}\in{U_{s}}$,
@@
Here and in what follows, $\mathfrak{v}\in 2^{\mathfrak{u}}$ says that $\mathfrak{v}$ is an element of the power set of $\mathfrak{u}$, which is equivalent to $\mathfrak{v}\subset\mathfrak{u}$.